Members of the Ski Club contributed equally to obtain $1800 for a holiday trip. When 6 members found that they could not go, their contributions were refunded and each remaining member then had to pay $10 more to raise the $1800. How many went on the trip?
Answer by checkley71(8403)

1800=m*c m=members & c=contribution
m=1800/c also
1800=(m-6)(c+10)
1800=(1800/c-6)(c+10)
1800=1800+18000/c-6c-60
-18000/c+60=-6c now multiply every term by c to eliminate the c in the denominator
-18000+60c=-6c^2
6c^2+60c-18000=0 divide all terms by 6
c^2+10c-3000=0
(c-50)(c+60)=0
c=50 & c=-60 the original cost is $50 then there were 1800/50=36 contributors
when 6 backed out there were 30 & they paid $60 each
proof
1800=30*60
1800=1800